OpenStudy (anonymous):

Multiply (z - 5)(z - 1)

OpenStudy (chrisasl):

z^2 - z - 5*z + 5 = z^2 - 6*z + 5 Would you mind me asking, if you are in middle school?

OpenStudy (anonymous):

Na im a senior this is math models

OpenStudy (anonymous):

is just that i suck at math

OpenStudy (anonymous):

Thanks for the help tho

OpenStudy (chrisasl):

Senior is the last year in high school? I asked that, in order to see how could I help you find it yourself :) If you need some more explaining tell me to draw something to help you

OpenStudy (anonymous):

Thanks chrisasl :)

OpenStudy (chrisasl):

 In your case (with the "-" ) we get this: \[(x - a) (x - b) = x^{2} - (a+b)*x +a*b\]